OpenStudy (queenbee567):

Help me with this equation : with solution set 1. Y/y-2 - 6/3y-1 < 1 (less than or equal to 1)

OpenStudy (phi):

is this \[ \frac{y}{y-2} - \frac{6}{3y-1} \le 1 \]?

OpenStudy (aakashtomar):

Yeah I think so. No way there's another variable 'Y' .

OpenStudy (phi):

the general idea is to compare to zero, so write the equation as \[ \frac{y}{y-2} - \frac{6}{3y-1} -1\le 0 \] and put everything over a common denominator.

OpenStudy (phi):

you then factor the left-hand side, and look for combinations of factors that give a negative product i.e. that make the left side less than (or equal) to zero.